Grade 12Mechanics

A body executes SHM on a line of length 16 cm with a frequency 1/24 find average velocity when it moves from mean position to half of amplitude???

Indraja
8 Years agoGrade 12

2 Answers

Eshan
8 Years ago
Dear student,

Let the SHM be represented asx=A sin\omega t
\implies v=A\omega cos\omega t

Hence average velocity overt=0tot=\dfrac{\pi}{6\omega}is
\dfrac{\int_0^{\pi/6\omega}A\omega cos\omega t}{\pi/6\omega}

=\dfrac{3A\omega}{\pi}

\omega=2\pi f=\dfrac{\pi}{12}
\implies v_{avg}=4cm/s
Khimraj
8 Years ago
time taken by body to reach half of amplitude from mean position T’ = T/12 
v = Acoswt
vavg = (1/T’)\int_{0}^{T`} Acos(wt)dt
vavg = (12A/Tw)[sinwt]_{0}^{T/12}
vavg = (6A/\pi)sin\pi/6
vavg = 3A/\pi
here A = 8cm
SO vavg = 24/\pi cm/s.
Hope it clears.